# Deep Reinforcement Learning for Real-Time Ground Delay Program Revision and Corresponding Flight Delay Assignments

### v1

@@ I Introduction
Over the past four decades, U.S. air traffic demand has increased more than sixfold, with an anticipated average growth rate of 2.2% over the next twenty years [1]; [2]. This escalating demand places additional strain on National Airspace System (NAS) components such as airports and sectors, and increases the workload for traffic controllers and managers. U.S. Air Traffic Management (ATM) encompasses Air Traffic Control (ATC) and the Traffic Flow Management System (TFMS). ATC coordinates the separation of aircraft through controller and pilot collaboration, while traffic flow managers employ various Traffic Management Initiatives (TMIs) to align traffic demand with capacity, strategically relocating delays to locations where they can be absorbed safely and efficiently. These TMIs, which have evolved significantly over decades, include diverse concepts, procedures, automation, and control characteristics.
@@
In practice, GDP design is largely manual, aided by the Flight Schedule Monitor, a decision-support tool [3];[4]. Traffic managers set the scope, duration, and program rate—reflecting airport capacity—which are then entered into the monitor to calculate ground delays for affected flights according to scheduling principles. Additionaly, GDP, a strategic Traffic Management Initiative (TMI), requires planning several hours in advance based on anticipated capacity and demand. However, these forecasts often fail to capture system uncertainties, leading to potential GDP revisions during operation if conditions change. This results in multiple GDP events, such as new issuances, revisions, or cancellations.
Improvement opportunities arise primarily from two types of uncertainties: supply (e.g., GDP program rate and time window) and demand (e.g., flight time variations and airline operational changes like cancellations and substitutions). There is a fundamental trade-off between control accuracy and the reliability of information over time. Initiating GDP earlier allows more flights to receive ground delays but increases the uncertainty in capacity and flight times. Despite these complexities, current real-time GDP designs remain suboptimal and require further enhancement [5]. Limited number of research focuses on optimizing GDP design and execution. Objective functions usually involve the weighted sum of ground and air delay, and airline equity with constraints from uncertainties regarding flight delays, weather conditions and airport capacity [5]; [6]; [7],[3]. Various solvers such as common machine learning, dynamic programming, stochastic optimization, and reinforcement learning, were applied to find optimal GDP parameters and to predict flight performance [8] [9] [10]. Due to the great complexity in the NAS and technological limitations, such models have yet to reach the stage of practical application.
@@ IV Conclusions
  • (3)
    Sensitivity Analysis: To better understand the influence of various operational parameters, future studies could modify the values assigned to different types of delays and congestion penalties. A systematic sensitivity analysis could help identify which factors most significantly affect GDP performance and could lead to more targeted improvements in model design.
  • (4)
    Uncertainty Handling: The next phase of model development should also focus on better handling of uncertainties associated with various stages of flight operations, such as gate departure times and taxi durations. Incorporating these uncertainties into the environmental inputs is essential for developing a more robust and reliable model that can effectively simulate and manage real-world conditions.
In conclusion, while our initial models provide a foundation for using RL in air traffic management, significant enhancements are required to fully harness their potential. By addressing the outlined limitations and incorporating more comprehensive data and sophisticated modeling techniques, future research can advance the state of the art in managing airport ground delays and overall air traffic efficiency.

### v2

@@ I Introduction
Over the past four decades, U.S. air traffic demand has increased more than sixfold, with an anticipated average growth rate of 2.2% over the next twenty years [1]; [2]; [3]. This escalating demand places additional strain on National Airspace System (NAS) components such as airports and sectors, and increases the workload for traffic controllers and managers. U.S. Air Traffic Management (ATM) encompasses Air Traffic Control (ATC) and the Traffic Flow Management System (TFMS) [4]. ATC coordinates the separation of aircraft through controller and pilot collaboration, while traffic flow managers employ various Traffic Management Initiatives (TMIs) to align traffic demand with capacity, strategically relocating delays to locations where they can be absorbed safely and efficiently. These TMIs, which have evolved significantly over decades, include diverse concepts, procedures, automation, and control characteristics.
@@
In practice, GDP design is largely manual, aided by the Flight Schedule Monitor, a decision-support tool [5];[6];[7]. Traffic managers set the scope, duration, and program rate—reflecting airport capacity—which are then entered into the monitor to calculate ground delays for affected flights according to scheduling principles. Additionaly, GDP, a strategic Traffic Management Initiative (TMI), requires planning several hours in advance based on anticipated capacity and demand. However, these forecasts often fail to capture system uncertainties, leading to potential GDP revisions during operation if conditions change. This results in multiple GDP events, such as new issuance, revisions, or cancellations.
Improvement opportunities arise primarily from two types of uncertainties: supply (e.g., GDP program rate and time window) and demand (e.g., flight time variations and airline operational changes like cancellations and substitutions). There is a fundamental trade-off between control accuracy and the reliability of information over time. Initiating GDP earlier allows more flights to receive ground delays but increases the uncertainty in capacity and flight times. Despite these complexities, current real-time GDP designs remain suboptimal and require further enhancement [8], [9]. Limited number of research focuses on optimizing GDP design and execution. Objective functions usually involve the weighted sum of ground and air delay, and airline equity with constraints from uncertainties regarding flight delays, weather conditions and airport capacity [8]; [10]; [11],[5]. Various solvers such as common machine learning, dynamic programming [12, 13], stochastic optimization [14, 15], and reinforcement learning, were applied to find optimal GDP parameters and to predict flight performance [16] [17] [18]. Due to the great complexity in the NAS and technological limitations, such models have yet to reach the stage of practical application.